OpenStudy (anonymous):

integral of 1/( e^(t+y) ) dy...?

OpenStudy (anonymous):

equal integral of e^-(t+y) = e^(-t-y)

OpenStudy (anonymous):

u=(-t-y) du- -dy

OpenStudy (anonymous):

=-1/(e^(t+y))

OpenStudy (anonymous):

what would du be if u = -t-y..?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

is it 1-dy..?

OpenStudy (anonymous):

-dy

OpenStudy (anonymous):

t is constant so its derivative is 0

OpenStudy (anonymous):

oo gosh ofcourse..thanks you!

OpenStudy (anonymous):

e^(t+y) = e^t * e^y hint

OpenStudy (anonymous):

and because e^t is a constant you can take it out of the integral and work with e^y